a PROJECTILE is thrown at an angle 30° with initial velocity 50 cm/s .when it is at the highest point it explodes and breaks into ratio 1:3 .the smaller part falls down vertically .find the distance where the heavier part lands.

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From momentum conservation in x and y direcn. separetely we get,
In x direcn.
Mvcos30 = M/4*0 + (3M/4)v’
v’ = 2*v*root(3)/3 m/s
In y direcn.
Mvsin30 = + v0*3M/4
v0 = 2v/3
Resultant velocity of heavier mass =>
=> root((2v/3)^2 + (2*v*root(3)/3)^2)
=> root(4v^2/9 + 4v^2/3)
=> root(16v^2/9)
=> 4v/3 m/s
Distance travelled,
(4v/3)^2 = v^2 + 2gs
we get,
s = 7v^2/18g
  = 7*0.25/180
  = 9.72222 x 10^-3 meter.